OpenStudy (anonymous):

1. Find the 9th term in the expansion of 7, 14, 28, . . . .

OpenStudy (anonymous):

an=7(9-1)7?

OpenStudy (anonymous):

\[7,7\times 2,7\times 2^2,7\times 2^3...\]

OpenStudy (anonymous):

ninth term would be \(7\times 2^8\)

OpenStudy (anonymous):

1792